OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x − 6y = 3 7x − 5y = −11 a. −2 b.−3 c. x can be any number as there are infinitely many solutions to the system. d.There is no x value as there is no solution to the system

OpenStudy (anonymous):

you can use substitution method.

OpenStudy (anonymous):

3x − 6y = 3 7x − 5y = −11 Multiply first equation by −7, and second equation by 3, then add: −21x + 42y = −21 21x − 15y = −33 ---------------------- 0x + 27y = −54 y = −2